Corporate Finance – Chapter 27- Short Term Financial PlanningQ. 09: The Rasputin Brewery is considering using a public warehouse loan as part of its short-term financing. The firm will r
Answer #1

Interest expense on laond = 540000 * 9.6% = 51840

Warehouse fees = 540000 * 1% = 540000 * 1% = 5400

Total charges = Interest expense + Warehouse fees = 51840 + 5400 = 57240

Warehouse expenses are paid at the beginning of the loan period and hence the net proceeds will be reduced by the warehouse fees.

Net proceeds = 540000 - 5400 = 534600

Effective annual rate = Total charges / Net proceeds = 57240 / 534600

= 10.71%